# EvRWKV: A Continuous Interactive RWKV Framework for Effective Event-Guided Low-Light Image Enhancement

### v2

@@ IV-A2 Implementation Details
All experiments were conducted with the Adam optimizer [21]. Learning rates were set to 1e41e{-4} for SDE datasets, 1.5e41.5e{-4} for SDSD datasets, and 1e41e{-4} for RELED datasets. The framework was trained on an NVIDIA RTX 3090 GPU for 80 training cycles with a batch size of eight. For the training set, we employed data-augmentation techniques consisting of random $256\times 256$ crops, horizontal flips, and rotations of $90^{\circ}$, $180^{\circ}$, and $270^{\circ}$. For the test set, we applied only a center crop of $256\times 256$.
@@ IV-A3 State-of-the-Art Methods for Comparisons
We compared our EvRWKV with ten methods, including five methods that only use RGB images as input(SNR-Net [55], Uformer [51], LLFlow-L-SKF [53], RetinexMac [33], Retinexformer [3]), five methods that use a RGB image and paired events as inputs(ELIE [17], eSL-Net [46], LLVE-SEG [35], ELEDNet [20], EvLight [30]). Since the source code of LLVE-SEG [35] is not publicly available, we implemented it according to the description in its paper. For the SNR-Net [55], Uformer [51], LLFlow-L-SKF [53], RetinexMac [33], Retinexformer [3], ELIE [17], eSL-Net [46], ELEDNet [20] and EvLight [30] methonds, we used the codes released by the authors to output their results.
IV-B Quantitative and Qualitative Evaluation
We evaluate the proposed EvRWKV framework on three challenging datasets—SDE, SDSD, and RELED—covering diverse LLIE scenarios. SDE and RELED contain real-world event-image pairs captured under indoor/outdoor and low-light blurry conditions respectively, while SDSD consists of synthetic events generated from dynamic video sequences. We use PSNR and SSIM as quantitative metrics and provide detailed qualitative analysis to validate the effectiveness of our approach.
@@ IV-B1 Evaluation on the SDE Dataset
For LLIE on the SDE dataset, the ability to simultaneously suppress noise and preserve fine structural details is the key evaluation criterion. We first qualitatively compare our EvRWKV method with several representative approaches including eSL-Net [46], ELEDNet [20], EvLight [30], and RetinexMac [33]. As shown in Fig. 6, RetinexMac and ELEDNet reduce noise to some extent but tend to oversmooth textures, resulting in loss of important details in dark regions. As shown in Fig. 6, EvLight maintains more structural information but suffers from residual noise and blur artifacts. eSL-Net, with a relatively small model size, provides limited enhancement capabilities and often underperforms in challenging scenes. In contrast, our EvRWKV method achieves a better balance between noise suppression and texture preservation, producing cleaner, brighter images with rich details and minimal artifacts. This demonstrates the effectiveness of our dual-domain fusion and adaptive gating mechanisms in handling complex low-light scenarios.
In terms of quantitative evaluation, Table I lists the PSNR and SSIM scores of different methods on the SDE dataset under both SDE-in and SDE-out settings. Our method achieves the highest scores of 23.09 PSNR and 0.766 SSIM for SDE-in, as well as 23.43 PSNR and 0.767 SSIM for SDE-out. These results outperform all compared methods including SNR-Net [55], Uformer [51], LLFlow-L-SKF [53], LLVE-SEG [35], Retinexformer [3], ELIE [17], LLVE-SEG [35], and the aforementioned qualitative comparison models. The significant improvements in both perceptual quality and objective metrics validate the superiority of our EvRWKV framework in LLIE.
@@ IV-B2 Evaluation on the SDSD Dataset
We further qualitatively compare the performance of different methods on the SDSD dataset for LLIE under challenging lighting and motion conditions. In Fig. 7, we observe that most methods improve the visibility and brightness of degraded images. However, RetinexMac [33], ELEDNet [20], and EvLight [30] exhibit limitations: As shown in Fig. 7, RetinexMac and ELEDNet tend to over-smooth textures and lose fine details, While EvLight, although preserving more structural information, causes noticeable color distortion in some regions. As shown in Fig. 7, eSL-Net [46] shows relatively limited enhancement capability due to its smaller model size. In comparison, our EvRWKV method delivers superior noise suppression and detail preservation, effectively enhancing both static and dynamic scene elements without introducing artifacts. The enhanced images show clearer textures, better contrast, and natural brightness, demonstrating the robustness of our dual-domain cross-modal fusion strategy.
Table II presents the PSNR and SSIM scores of different methods evaluated on the SDSD dataset. Our method achieves the highest scores in both SDSD-in and SDSD-out scenarios, with PSNR values of 28.96 and 27.93, and SSIM values of 0.920 and 0.839, respectively. These results outperform all other compared methods, validating the strong restoration ability and noise robustness of our EvRWKV framework under various low-light conditions. Overall, our approach demonstrates both qualitative and quantitative advantages over existing state-of-the-art techniques on the SDSD dataset.
@@ IV-B3 Evaluation on the RELED Dataset
Figure 8: Visual comparisons with state-of-the-art methods on the Reled dataset.
TABLE III: Ablation study of key components on the SDSD-in dataset.
@@
w/o EISFE 27.35 0.894
w/o SpatialMix 26.88 0.798
w/o ChannelMix 26.60 0.896
Ours 28.96 0.920
To comprehensively evaluate the enhancement performance of our EvRWKV method on the RELED dataset, we compare it qualitatively with several representative approaches including eSL-Net [46], ELEDNet [20], EvLight [30], and RetinexMac [33]. As shown in Fig. 8, the RELED dataset contains extremely low-light images with severe motion blur and noise. As shown in Fig. 8, eSL-Net often causes overexposure while attempting to enhance brightness. RetinexMac boosts overall brightness but fails to eliminate noticeable blur. ELEDNet and EvLight aim to balance illumination, but this often leads to the loss of finer details. In contrast, our EvRWKV method effectively leverages event-based motion cues and image textures to suppress noise, reduce blur, and restore fine details, producing visually clearer, sharper, and more natural images without color artifacts or overexposure.
Table I reports the PSNR and SSIM scores for different methods on the RELED dataset. Our EvRWKV achieves the best quantitative results with a PSNR of 32.18 and SSIM of 0.910, significantly outperforming other methods including eSL-Net, ELEDNet, EvLight, and RetinexMac. These quantitative gains demonstrate the robustness and generalization capability of our dual-domain cross-modal fusion framework in handling challenging low-light and motion-degraded scenarios. Overall, our method achieves excellent qualitative and quantitative results across all three real-world LLIE benchmarks.
@@ IV-C Ablation Study
We conduct comprehensive ablation studies on the SDSD-in dataset to evaluate the contribution of each key component in our EvRWKV framework. In particular, we assess the effectiveness of the EISFE module, the SpatialMix and ChannelMix mechanisms within the Cross-RWKV module, and the role of each individual loss term in the overall objective function. For each study, we remove or modify a single module while keeping the rest of the framework unchanged to isolate its impact. The results clearly demonstrate the necessity and complementarity of these components in achieving robust and high-quality LLIE. As shown in Fig. 9, the removal of EISFE (b) results in color distortion. Without SpatialMix (c), visible noise and blurred textures appear, while omitting ChannelMix (d) leads to flat tonal rendering. The full model (e) demonstrates the best visual clarity and color fidelity, especially in complex shadow regions.
@@ IV-C1 Impact of EISFE Module
To assess the contribution of frequency-aware fusion, we remove the EISFE module and directly decode the features after the Cross-RWKV backbone. As shown in Table III, the absence of EISFE leads to a 1.61 dB drop in PSNR and 0.026 reduction in SSIM. This indicates the importance of frequency-domain filtering and spatial deformable convolution in denoising and structure restoration. Visually, the model without EISFE shows residual noise and blurred edges.
@@ IV-C2 Impact of SpatialMix and ChannelMix
To verify the effectiveness of the spatial and channel interaction designs in Cross-RWKV, we conduct ablation studies by individually removing each module. As shown in Table III, removing the SpatialMix results in a 2.08 dB PSNR drop and 0.122 SSIM decline, highlighting its critical role in capturing long-range dependencies and facilitating event-image alignment. On the other hand, removing ChannelMix leads to a 2.36 dB decrease in PSNR and a 0.024 SSIM reduction, indicating its importance in refining semantic consistency across feature channels. These results confirm that both modules contribute distinctly and significantly to performance. As illustrated in Fig. 9, ChannelMix aids in preserving local contrast, while SpatialMix enhances spatial structure and reduces texture inconsistency.
@@ IV-C3 Impact of Loss Functions
To evaluate the contribution of each loss term, we progressively remove pixel, perceptual, SSIM, and MS-SSIM losses from the training objective. The results are summarized in Table IV. Removing the reconstruction loss $L_{r}$ results in the lowest performance, with a PSNR of 26.84 and an SSIM of 0.896. Excluding the perceptual loss $L_{p}$ leads to a slight reduction in performance, with a PSNR of 26.18 and an SSIM of 0.885. Removing the SSIM loss $L_{s}$ results in a PSNR of 26.06 and an SSIM of 0.876, reflecting a loss in structural preservation. Finally, excluding the multi-scale SSIM loss $L_{m}$ leads to a PSNR of 27.84 and an SSIM of 0.899. The full model, with all losses included, achieves the best results with a PSNR of 28.96 and an SSIM of 0.920, demonstrating a 0.84 dB gain in PSNR over the setting without $L_{r}$. This confirms the complementary benefits of each loss term in guiding high-quality reconstruction.
TABLE IV: Ablation study of loss functions on the SDSD-in dataset.
@@
w/o reconstruction loss ($L_{r}$) 26.84 0.896
w/o perceptual loss ($L_{p}$) 26.18 0.885
w/o SSIM loss ($L_{s}$) 26.06 0.876
w/o MS-SSIM loss ($L_{m}$) 27.84 0.899
Ours 28.96 0.920
IV-D Limitations
While EvRWKV performs well across diverse low-light conditions, it encounters challenges in extremely dark scenes with sparse event activity and limited structural cues. In such cases, the framework may struggle to recover fine details or accurate colors, occasionally leading to over-smoothing or artifacts. Additionally, severe motion blur or sensor noise can impair spatial-temporal alignment. Future work will focus on improving robustness through uncertainty modeling and adaptive fusion strategies for broader generalization.

### v3

@@ IV-A2 Implementation Details
All experiments were conducted with the Adam optimizer [kingma2014adam]. Learning rates were set to 1e41e{-4} for SDE datasets, 1.5e41.5e{-4} for SDSD datasets, and 1e41e{-4} for RELED datasets. The framework was trained on an NVIDIA RTX 3090 GPU for 80 training cycles with a batch size of eight. For the training set, we employed data-augmentation techniques consisting of random $256\times 256$ crops, horizontal flips, and rotations of $90^{\circ}$, $180^{\circ}$, and $270^{\circ}$. For the test set, we applied only a center crop of $256\times 256$.
@@ IV-A3 State-of-the-Art Methods for Comparisons
We compared our EvRWKV with ten methods, including five methods that only use RGB images as input(SNR-Net [xu2022snr], Uformer [wang2022uformer], LLFlow-L-SKF [wu2023learning], RetinexMac [10810469], Retinexformer [cai2023retinexformer]), five methods that use a RGB image and paired events as inputs(ELIE [jiang2023event], eSL-Net [wang2020event], LLVE-SEG [liu2023low], ELEDNet [kim2024towards], EvLight [liang2024towards]). Since the source code of LLVE-SEG [liu2023low] is not publicly available, we implemented it according to the description in its paper. For the SNR-Net [xu2022snr], Uformer [wang2022uformer], LLFlow-L-SKF [wu2023learning], RetinexMac [10810469], Retinexformer [cai2023retinexformer], ELIE [jiang2023event], eSL-Net [wang2020event], ELEDNet [kim2024towards] and EvLight [liang2024towards] methonds, we used the codes released by the authors to output their results.
IV-B Comparison and Evaluation
We evaluate the proposed EvRWKV framework on three challenging datasets, namely SDE, SDSD, and RELED, covering diverse LLIE scenarios. SDE and RELED contain real-world event-image pairs captured under indoor and outdoor low-light conditions as well as motion blur scenarios, while SDSD consists of synthetic events generated from dynamic video sequences. We use PSNR, SSIM, LPIPS, and NIQE as quantitative metrics to assess both pixel-level accuracy and perceptual quality, and provide detailed qualitative analysis to validate the effectiveness of our approach.
@@ IV-B1 Evaluation on the SDE Dataset
For LLIE on the SDE dataset, the ability to simultaneously suppress noise and preserve fine structural details is the key evaluation criterion. We first qualitatively compare our EvRWKV method with several representative approaches including eSL-Net [wang2020event], ELEDNet [kim2024towards], EvLight [liang2024towards], and RetinexMac [10810469]. As shown in Fig. 7, RetinexMac and ELEDNet reduce noise to some extent but tend to oversmooth textures, resulting in loss of important details in dark regions. As shown in Fig. 7, EvLight maintains more structural information but suffers from residual noise and blur artifacts. eSL-Net, with a relatively small model size, provides limited enhancement capabilities and often underperforms in challenging scenes. In contrast, our EvRWKV method achieves a better balance between noise suppression and texture preservation, producing cleaner, brighter images with rich details and minimal artifacts. This demonstrates the effectiveness of our dual-domain fusion and adaptive gating mechanisms in handling complex low-light scenarios.
In terms of quantitative evaluation, Table I presents comprehensive metrics for different methods on the SDE dataset. Our method achieves the highest PSNR of 23.09 and SSIM of 0.770 for SDE-in, along with the best LPIPS of 0.094 and NIQE of 8.534, demonstrating superior performance in both pixel-level reconstruction and perceptual quality. For SDE-out, our method obtains 23.43 PSNR, 0.768 SSIM, 0.091 LPIPS, and 8.926 NIQE, consistently outperforming all compared methods including SNR-Net [xu2022snr], Uformer [wang2022uformer], LLFlow [wu2023learning], Retinexformer [cai2023retinexformer], ELIE [jiang2023event], EvLight [liang2024towards], and ELEDNet [kim2024towards]. The significant improvements across both traditional metrics and perceptual metrics validate the superiority of our EvRWKV framework in LLIE.
@@ IV-B2 Evaluation on the SDSD Dataset
We further qualitatively compare the performance of different methods on the SDSD dataset for LLIE under challenging lighting and motion conditions. In Fig. 8, we observe that most methods improve the visibility and brightness of degraded images. However, RetinexMac [10810469], ELEDNet [kim2024towards], and EvLight [liang2024towards] exhibit limitations: RetinexMac and ELEDNet tend to over-smooth textures and lose fine details, While EvLight, although preserving more structural information, causes noticeable color distortion in some regions. As shown in Fig. 8, eSL-Net [wang2020event] shows relatively limited enhancement capability due to its smaller model size. In comparison, our EvRWKV method delivers superior noise suppression and detail preservation, effectively enhancing both static and dynamic scene elements without introducing artifacts. The enhanced images show clearer textures, better contrast, and natural brightness, demonstrating the robustness of our dual-domain cross-modal fusion strategy.
Table II presents comprehensive quantitative evaluation on the SDSD dataset. Our method achieves the best performance across multiple metrics in both SDSD-in and SDSD-out scenarios. Specifically, for SDSD-in, we obtain PSNR of 28.96, SSIM of 0.920, LPIPS of 0.032, and NIQE of 6.451, while for SDSD-out, we achieve 27.93 PSNR, 0.839 SSIM, 0.058 LPIPS, and 4.110 NIQE. These results consistently outperform all compared methods, demonstrating superior performance in both pixel-level accuracy and perceptual quality. The improvements in LPIPS and NIQE further validate that our enhanced images better align with human visual perception and exhibit more natural image characteristics. Overall, our approach demonstrates both qualitative and quantitative advantages over existing state-of-the-art techniques on the SDSD dataset.
@@ IV-B3 Evaluation on the RELED Dataset
To comprehensively evaluate the enhancement performance of our EvRWKV method on the RELED dataset, we compare it qualitatively with several representative approaches including eSL-Net [wang2020event], ELEDNet [kim2024towards], EvLight [liang2024towards], and RetinexMac [10810469]. As shown in Fig. 9, the RELED dataset contains extremely low-light images with severe motion blur and noise. As shown in Fig. 9, eSL-Net often causes overexposure while attempting to enhance brightness. RetinexMac boosts overall brightness but fails to eliminate noticeable blur. ELEDNet and EvLight aim to balance illumination, but this often leads to the loss of finer details. In contrast, our EvRWKV method effectively leverages event-based motion cues and image textures to suppress noise, reduce blur, and restore fine details, producing visually clearer, sharper, and more natural images without color artifacts or overexposure.
Table II reports comprehensive metrics for different methods on the RELED dataset. Our EvRWKV achieves the best quantitative results with PSNR of 32.18, SSIM of 0.911, LPIPS of 0.044, and NIQE of 9.207, significantly outperforming other methods including eSL-Net [wang2020event], ELEDNet [kim2024towards], EvLight [liang2024towards], and RetinexMac [10810469]. These comprehensive quantitative gains across both traditional and perceptual metrics demonstrate the robustness and generalization capability of our dual-domain cross-modal fusion framework in handling challenging low-light and motion-degraded scenarios. Overall, our method achieves excellent qualitative and quantitative results across all three real-world LLIE benchmarks.
TABLE IV: Ablation study of key components on the SDSD-in dataset.
@@
27.35 0.894
26.88 0.798
26.60 0.896
28.96 0.920
TABLE V: Ablation study of loss functions on the SDSD-in dataset.
[…]
0 0.8 1 0.5 26.84 0.896
1 0 1 0.5 26.18 0.885
1 0.8 0 0.5 26.06 0.876
1 0.8 1 0 27.84 0.905
1 0.8 1.5 0.5 27.98 0.908
1 0.8 1.5 1 26.12 0.891
1 1 1 1 26.73 0.891
1 0.8 1 0.5 28.96 0.920
TABLE VI: Ablation study of voxel bin size on the RELED dataset.
[…]
$B=16$ 30.42 0.879
$B=64$ 30.57 0.881
$B=32$ (Ours) 32.18 0.911
TABLE VII: Ablation study of Cross-RWKV depth on the SDSD-in dataset.
[…]
3 [1, 2, 2] 25.63 0.891 4.3
3 [2, 4, 4] 26.04 0.896 6.6
5 [1, 2, 2, 4, 4] 29.02 0.911 82.2
5 [2, 4, 4, 8, 8] 27.60 0.898 147
4 [2, 4, 4, 8] 28.18 0.901 33.1
4 (Ours) [1, 2, 2, 6] 28.96 0.920 24.2
IV-B4 Generalization Comparison
To evaluate the generalization capability and analyze the limitations of our method, we conduct cross-dataset experiments using the model trained on SDE-out and directly testing it on the LIE dataset [jiang2023event] without any fine-tuning. Table III shows that our method achieves the best performance with PSNR of 20.75 and SSIM of 0.645. As shown in Fig. 10, our method effectively recovers details in low-light images, while simultaneously preventing overexposure, as seen in (e), and color distortion, as seen in (d). These results demonstrate the strong cross-dataset generalization capability and robustness of EvRWKV.
Figure 11: Representative failure cases in challenging scenarios.
@@ IV-C Ablation Study
We conduct comprehensive ablation studies to evaluate the contribution of each key component and hyper-parameter in our EvRWKV framework. We assess the effectiveness of the EISFE module, the SpatialMix and ChannelMix mechanisms within the Cross-RWKV Module, the role of each loss term, as well as the influence of critical hyper-parameters including Cross-RWKV depth and voxel bin size. For each study, we isolate the impact by removing or modifying a single component while keeping the rest unchanged. The results clearly demonstrate the necessity and complementarity of these components in achieving robust and high-quality LLIE. As shown in Fig. 12, the removal of EISFE (b) results in color distortion. Without SpatialMix (c), visible noise and blurred textures appear, while omitting ChannelMix (d) leads to flat tonal rendering. The full model (e) demonstrates the best visual clarity and color fidelity, especially in complex shadow regions.
@@ IV-C1 Impact of EISFE Module
To assess the contribution of frequency-aware fusion, we remove the EISFE module and directly decode the features after the Cross-RWKV backbone. As shown in Table IV, the absence of EISFE leads to a 1.61 dB drop in PSNR and 0.026 reduction in SSIM. This indicates the importance of frequency-domain filtering and spatial deformable convolution in denoising and structure restoration. Visually, the model without EISFE shows residual noise and blurred edges.
@@ IV-C2 Impact of SpatialMix and ChannelMix
To verify the effectiveness of the spatial and channel interaction designs in Cross-RWKV, we conduct ablation studies by individually removing each module. As shown in Table III, removing the SpatialMix results in a 2.08 dB PSNR drop and 0.122 SSIM decline, highlighting its critical role in capturing long-range dependencies and facilitating event-image alignment. On the other hand, removing ChannelMix leads to a 2.36 dB decrease in PSNR and a 0.024 reduction in SSIM, indicating its importance in refining semantic consistency across feature channels. These results confirm that both modules contribute distinctly and significantly to performance. As illustrated in Fig. 12, ChannelMix aids in preserving local contrast, while SpatialMix enhances spatial structure and reduces texture inconsistency.
@@ IV-C3 Impact of Loss Functions
To validate our loss configuration, we conducted ablation studies by both removing individual loss terms and adjusting their weights, as detailed in Table V. Removing either the SSIM loss $L_{s}$ or the perceptual loss $L_{p}$ causes a significant performance drop to 26.06 dB and 26.18 dB respectively, highlighting their crucial role in preserving structural and perceptual quality. While other configurations with different weightings show competitive results, our final empirically determined weights yield the optimal performance of 28.96 dB PSNR and 0.920 SSIM. This confirms that both the selection and weighting of the loss terms are critical for optimal results.
IV-C4 Impact of Voxel Bin Size
The voxel bin size $B$ determines the temporal resolution of event representation. We evaluate three settings on the RELED dataset. As shown in Table VI, $B=16$ yields a PSNR of 30.42 dB, while $B=64$ achieves 30.57 dB. Our default setting $B=32$ achieves the best performance with 32.18 dB PSNR and 0.911 SSIM. This indicates that $B=32$ provides an optimal balance: smaller bins suffer from excessive sparsity, while larger bins lose temporal precision.
IV-C5 Impact of Cross-RWKV Depth
The number of stacked Cross-RWKV blocks affects the model’s capacity to capture long-range dependencies. As shown in Table VII, using 3 levels results in 25.63 dB PSNR with only 4.3M parameters, indicating insufficient capacity. Increasing to 5 levels achieves 29.02 dB but requires 82.2M parameters. Our 4-level configuration achieves 28.96 dB PSNR and 0.920 SSIM with 24.2M parameters, providing the best balance between performance and computational cost.
Figure 13: Qualitative comparison of semantic segmentation results for low-light scenes. The baseline method (i.e., SAM [kirillov2023segment]) uses the raw low-light image as input, SAM+EvLight uses the EvLight-enhanced image as input, while SAM+Ours uses our enhanced image as input. GT shows the segmentation on GT images.
TABLE VIII: Quantitative results of semantic segmentation on our SDE-in and RELED dataset.
Method SDE-in RELED
@@ IV-C5 Impact of Cross-RWKV Depth
SAM [kirillov2023segment] 60.46 30.84 40.81 68.64 20.13 29.13
SAM [kirillov2023segment]+EvLight 66.38 38.16 49.52 78.26 37.27 46.70
SAM [kirillov2023segment]+Ours 69.27 41.77 52.55 80.90 39.37 48.25
Improve (%) +4.35 +9.46 +6.12 +3.37 +5.63 +3.32
IV-D Downstream Applications
While event cameras offer unique advantages for low-light perception, leveraging their complementary information to enhance frame quality enables the direct application of well-established frame-based vision algorithms without requiring specialized event-based processing. To demonstrate the practical utility of our enhancement method, we conduct semantic segmentation experiments using the off-the-shelf ViT-H SAM model [kirillov2023segment] on the SDE-in and RELED datasets. We compare segmentation performance across four types of inputs: raw low-light images, EvLight-enhanced images [liang2024towards], our enhanced images, and ground-truth (GT). Since pixel-level segmentation annotations are not available, we use SAM segmentation results on GT as pseudo GT, which provides a reliable reference given SAM’s optimal performance on high-quality, well-exposed images. For quantitative evaluation, we report average pixel accuracy (aAcc), mean Intersection over Union (mIoU), and mean class accuracy (mAcc) in Table VIII. As shown in Fig. 13, our enhancement better preserves discriminative features in challenging regions with fine-grained structures compared to both raw images and EvLight, producing segmentation masks with greater consistency to the GT. These results confirm that our EvRWKV effectively recovers semantic content essential for practical applications such as autonomous driving and robotic navigation in low-light environments.
IV-E Limitations
Despite the strong performance, EvRWKV encounters challenges in certain extreme scenarios, as illustrated in Fig. 11. First, in extremely dark scenes where event activity becomes highly sparse due to minimal brightness changes, our method struggles to extract sufficient motion cues for effective cross-modal fusion. As shown in Fig. 11(a), the scarcity of events in such conditions prevents our Cross-RWKV Module from establishing reliable correspondences with image features, occasionally leading to over-smoothing or color artifacts. Second, when RGB frames suffer from severe motion blur combined with sensor noise, both modalities are simultaneously degraded. As illustrated in Fig. 11(b), since event voxels aggregate multiple temporal instances into a single frame representation, critical motion information is lost during this temporal compression, limiting our method’s ability to recover sharp details from blurred inputs. Third, in cross-dataset scenarios with rapid illumination changes, as shown in Fig. 11(c), sudden lighting transitions trigger bursts of events that do not correspond to actual scene structure, interfering with our spatial-temporal alignment and resulting in degraded outputs. These observations motivate future work on more robust event processing and domain adaptation techniques to overcome these extreme scenarios.